Checkliste für Fachschaftsfahrten etc.

Paul Ebermann 11. Mai 2004

Inhaltsverzeichnis

1 Privates 1

1 Privates

• Waschzeug – Handtuch

– Zahnbürste

– Zahnpasta

– Kamm

• Hausschuhe

• Kleidung (genug)

• Badehose

• Trinkflasche

• Schlafanzug

• Ohrstöpsel